OpenStudy (anonymous):

Find the equation of the tangent line to the graph of y=9^x at x=1.

OpenStudy (whpalmer4):

Are you taking calculus?

OpenStudy (anonymous):

In a general sense yes.

OpenStudy (whpalmer4):

Okay, I ask because I'm going to propose that you take the first derivative of \(y = 9^x\), and if you don't know calculus, you're going to say "huh?" :-)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Oh yah, I know the derivative. \[y=9^xln(9)\]

OpenStudy (whpalmer4):

I like you already :-) The function is \(y = f(x) = 9^x\) and the derivative is \(y' = f'(x) = 9^x\ln(9)\) so we can evaluate the derivative at \(x=1\) to find the slope of the tangent line. Then we evaluate the function \(f(1)\) to find the \((x_0,y_0)\) that we need to fit the line through. We fit the tangent line with the point-slope equation: \[y - y_0 = m(x-x_0)\]where \(m\) is the value of \(f'(1) = 9^1\ln(9)\)

OpenStudy (anonymous):

so \[m \approx19.78\] then just plug that into \[y-y_0=m(x-x_0)\]

OpenStudy (whpalmer4):

When we do that, we get a nice tangent line...